# SinLlama - A Large Language Model for Sinhala

### v1

@@ Abstract
Low-resource languages such as Sinhala are often overlooked by open-source Large Language Models (LLMs). In this research, we extend an existing multilingual LLM (Llama-3-8B) to better serve Sinhala. We enhance the LLM tokenizer with Sinhala specific vocabulary and perform continual pre-training on a cleaned 10 million Sinhala corpus, resulting in the SinLlama model. This is the very first decoder-based open-source LLM with explicit Sinhala support. When SinLlama was instruction fine-tuned for three text classification tasks, it outperformed base and instruct variants of Llama-3-8B by a significant margin.
@@ I Introduction
Large Language Models (LLMs) such as ChatGPT have significantly advanced Natural Language Processing (NLP) for languages with abundant training data and linguistic resources. While some recent successful open-source LLMs such as Llama [14] and Gemma [32] claim to have multilingual support, such multilingual LLMs (multiLLMs) continue to under-perform on low-resource languages (LRLs) [41]. Given that these LLMs are being created by large tech corporations, it is fair to assume that the decision on which languages to include in an LLM is driven by the perceived economic benefits in a global scale. As such, there is a risk of LRLs continue to be neglected, thus further increasing the digital divide between communities.
Sinhala is an Indo-Aryan language, which is only being used by a population of about 20 million in the island nation of Sri Lanka [10]. According to Ranathunga and De Silva [29]’s language categorization that uses the category definition of Joshi et al. [18], Sinhala is a low-resource language. There has been previous research that resulted in word embeddings and encoder-based language models for Sinhala [11]. However, techniques based on such models have become out-dated due to the demonstrated potential of LLMs. However, Sinhala, similar to many other LRLs, has not been considered during pre-training of any of the popular open-source LLMs such as Llama and Gemma released by the tech giants. While Sinhala is included in MaLA-500 [23] and Aya-101 [35] released by research teams, these models have not been able to outperform models such as Llama. On the other hand, MaLA-500 and Aya-101 models are large in size, incurring in large computer memory requirements.
To address these limitations, we present a focused adaptation of a state-of-the-art multiLLM to better support Sinhala. Our approach involves two key stages: (1) enhancing the model’s tokenizer by merging Sinhala-specific tokens into its vocabulary to improve subword segmentation; and (2) conducting continual pretraining using a 10 million unlabeled sentences corpus that includes cleaned and diverse Sinhala text, ensuring deeper linguistic representation.
We first carried out an investigation on the available open-source LLMs to select the best performing LLM for Sinhala. Llama-3 emerged as the most promising LLM. Then we further pre-trained the Llama-3-8B model as mentioned above. We term the resulting model, SinLlama. This model will be publicly released upon paper acceptance. In order to evaluate the performance of SinLlama against the original Llama-3, we fine-tuned both with data from three Sinhala benchmarks [37]: writing-style classification, news categorization and sentiment analysis. The results show that when fine-tuned on Sinhala task-specific data, SinLlama significantly outperforms both fine-tuned Llama-base and Llama-instruct versions.
@@ II-A Large Language Models (LLMs)
Pre-trained Language Models (PLMs), created using the Transformer architecture [36], have significantly advanced the NLP landscape. Among them, decoder-only models such as ChatGPT have become especially popular due to their strong performance. Open-source large-scale decoder models, known as LLMs (e.g., LLaMA [14], Mistral [17], Gemma [32]), are trained on much larger datasets and offer performance comparable to commercial systems. Training an LLM has two primary stages: training a tokenizer and self-supervised training with an unlabeled raw text corpus. The latter step is termed pre-training, and the resulting model is termed a base model. However, such base models lack task-specific capabilities. Therefore, these base models should be fine-tuned with task-specific instruction data, which is also termed instruct-tuning. The aforementioned open-source LLMs are available in both base and insturct versions.
@@ II-B Empirical Studies on LLM Performance for LRLs
The evaluation of LLMs across multiple languages reveals significant performance disparities between high-resource and LRLs. Ahuja et al. [2]’s MEGA framework conducted the first comprehensive benchmarking of LLMs on 16 NLP tasks across 70 diverse languages (including Sinhala) for comparing the performance of different models on XLSUM dataset. They showed that while models such as GPT-4 performed well in high-resource languages, they struggle with LRLs, especially those with unique morphological structures or scripts.
@@ II-D Sinhala Language Computing
While there has been research on implementing Sinhala language computing tools, most of has been done using rule-based, statistical or early-generation deep learning techniques [10]. In the NLP field, these techniques have long become obsolete. Nevertheless, there have been some attempts to keep Sinhala NLP up-to-date. Lakmal et al. [21] presented Word2Vec and FastText embeddings for Sinhala and carried out a detailed investigation. Dhananjaya et al. [11] built the encoder-based language model SinBERT, by fine-tuning an encoder-based Transformer [24] with 15.7 million Sinhala sentences.
@@ III Model Selection
In the initial phase of our research, we aimed to identify the most suitable open-source multiLLM, focusing on LRLs. The first round of filtration was done referring to empirical research that compared open-source LLM. This resulted in the following models: Llama-3-8B [14], MaLA-500 [23], Aya-101 [35], PolyLM [38], XGLM, [22] Falcon [3], BLOOMZ [26], Phi3 [1], Mistral [17], OPT [40], Aya-23 [35], and BigTrans [39]. These models are listed in Table I. Based on criteria including multilingual capability, parameter size (fewer than 8 billion parameters) [^1] Model size has a direct relationship with the memory requirement in using an LLM., Sinhala language representation, and performance metrics, we selected Llama-3-8B [14] for further experimentation.
@@
Model Parameters Official Sinhala Support No:of languages
@@ IV-A Pre-training Data
Pre-training of an LLM has to be done with a substantially large dataset, in order to guarantee it learns the language-specific information. Therefore, we created a dataset by combining MADLAD-400 [20] and CulturaX [27] for the Sinhala language [^2] The dataset created by Dhananjaya et al. [11] is no longer accessible. MADLAD-400 is a manually audited, general-domain 3 Trillion token monolingual dataset based on CommonCrawl [28], spanning 419 languages. CulturaX is a multilingual dataset with 6.3 Trillion tokens in 167 languages. Although both have been cleaned, there were some entries that contained non-Sinhala words. Because of that, heuristics-based filtering [13] was applied to remove such noisy sentences. These include rules such as minimum sentence length, non-Sinhala sentence removal, exclusion of sentences with irregular punctuation, and elimination of URLs and numeric prefixes.For the pre-training, we combined both datasets and removed duplicates. Our final dataset contains 10,730,154 (10.7M) Sinhala sentences, comprising 303,958,959 (303.96M) tokens. This dataset will be released upon paper acceptance.
@@ V Continual Pretraining to Build SinLlama
We used vocabulary extension and continuous pre-training (see Section II-C) on Llama-3-8B to create SinLlama. We trained a tokenizer using the pre-training corpus using tiktoken [^3] https://github.com/openai/tiktoken, and merged this tokenizer with the Llama 3 tokenizer. We used this extended tokenizer and performed continual pretraining using the codebase from Chinese-Llama [8]. The only modification made to the original Chinese-Llama hyperparameters was the reduction of the block_size from 1024 to 512444Larger block sizes cause memory allocation issues on the GPUs we used, preventing successful model loading..
@@ VI-A Prompt Selection
For both inference and supervised fine-tuning, it was essential to select the most effective prompt templates from the available datasets. We adopted the prompt format introduced by the Alpaca Instruct Template [5] as in Fig.  1. To identify the most suitable prompts for each classification task, we referred to the work by Ahuja et al. [2], which provided well-designed prompts for several tasks.
@@ A-A Writing-style-classification
Feature columns in the dataset are ”comments”, ”labels”.
@@ A-B Sentiment-tagger
The original dataset consist of XML tags. After the processing we created a dataset with the following features. Feature columns in the dataset are ”context”, ”text”, ”sentiment”.
@@ A-C Sinhala-News-Category-classification
Feature columns in the dataset are ”comments”, ”labels”.

### v2

@@ Abstract
Low-resource languages such as Sinhala are often overlooked by open-source Large Language Models (LLMs). Therefore, it is imperative that the existing LLMs are further trained to cover such languages. In this research, we extend an existing multilingual LLM (Llama-3-8B) to get a better coverage for Sinhala. We enhanced the LLM tokenizer with Sinhala specific vocabulary and performed continual pre-training on a 10 million sentence Sinhala corpus, resulting in the SinLlama model. This is the very first decoder-based open-source LLM with explicit Sinhala support. When SinLlama was instruction fine-tuned for three text classification tasks, it outperformed base and instruct variants of Llama-3-8B by a significant margin.
@@ I Introduction
Large Language Models (LLMs) such as ChatGPT have significantly advanced the field of Natural Language Processing (NLP) for languages with abundant training data and linguistic resources. However, their support for low-resource languages (LRLs) is significantly low. While some recent successful open-source LLMs such as Llama [1] and Gemma [2] claim to have multilingual support, such multilingual LLMs (multiLLMs) continue to under-perform on LRLs [3]. Given that these LLMs are being created by large tech corporations, it is fair to assume that the decision on which languages to include in an LLM is driven by the perceived economic benefits in a global scale. As such, there is a risk of LRLs continue to be neglected, thus further increasing the digital divide between communities.
Sinhala is an Indo-Aryan language, which is only being used by a population of about 20 million in the island nation of Sri Lanka [4]. According to Ranathunga and De Silva [5]’s language categorization that uses the category definition of Joshi et al. [6], Sinhala is a low-resource language. There has been previous research that resulted in word embeddings and encoder-based language models for Sinhala [7]. However, techniques based on such models have become out-dated due to the demonstrated potential of LLMs. However, Sinhala, similar to many other LRLs, has not been considered during pre-training of any of the popular open-source LLMs such as Llama and Gemma released by the tech giants. While Sinhala is included in MaLA-500 [8] and Aya-101 [9] released by research teams, these models have not been able to outperform models such as Llama. On the other hand, MaLA-500 and Aya-101 models are large in size, incurring in large computer memory requirements when used for downstream tasks.
To address these limitations, we present a focused adaptation of a state-of-the-art multiLLM to better support Sinhala. Our approach involves two key stages: (1) enhancing the model’s tokenizer by merging Sinhala-specific tokens into its vocabulary to improve subword segmentation; and (2) conducting continual pretraining using a 10 million sentence corpus that includes cleaned and diverse Sinhala text, ensuring deeper linguistic representation.
We first carried out an investigation on the available open-source LLMs to select the best performing LLM for Sinhala. Llama-3 emerged as the most promising LLM. Then we further pre-trained the Llama-3-8B model as mentioned above. We term the resulting model, SinLlama [^1] https://huggingface.co/polyglots/SinLlama_v01, which is publicly released. In order to evaluate the performance of SinLlama against the original Llama-3, we fine-tuned both with data from three Sinhala benchmark datasetss [10]: writing-style classification, news categorization and sentiment analysis. The results show that when fine-tuned on Sinhala task-specific data, SinLlama significantly outperforms both fine-tuned Llama-base and Llama-instruct versions.
@@ II-A Large Language Models (LLMs)
Pre-trained Language Models (PLMs), created using the Transformer architecture [11], have significantly advanced the NLP landscape. Among them, decoder-only models such as ChatGPT have become especially popular due to their strong performance across a multitude of tasks. Open-source large-scale decoder models, known as LLMs (e.g., LLaMA, Mistral [12], Gemma), are also trained on larger datasets and offer performance comparable to commercial systems. Training an LLM has two primary stages: training a tokenizer and self-supervised training with an unlabeled raw text corpus. The latter step is termed pre-training, and the resulting model is termed a base model. However, such base models lack task-specific capabilities. Therefore, these base models should be fine-tuned with task-specific instruction data, which is termed fine-tuning or instruct-tuning. The aforementioned open-source LLMs are available in both base and instruct versions.
@@ II-B Empirical Studies on LLM Performance for LRLs
The evaluation of LLMs across multiple languages reveals significant performance disparities between high-resource and LRLs. Ahuja et al. [13]’s MEGA framework conducted the first comprehensive benchmarking of LLMs on 16 NLP tasks across 70 diverse languages (including Sinhala) for comparing the performance of different models on XLSUM dataset. They showed that while models such as GPT-4 performed well in high-resource languages, they struggle with LRLs, especially those with unique morphological structures and/or scripts.
@@ II-D Sinhala Language Computing
While there has been research on implementing Sinhala language computing tools, most of it has been done using rule-based, statistical or early-generation deep learning techniques [4]. In the NLP field, these techniques have long become obsolete. Nevertheless, there have been some attempts to keep Sinhala NLP up-to-date. Lakmal et al. [20] presented Word2Vec and FastText embeddings for Sinhala and carried out a detailed investigation. Dhananjaya et al. [7] built the encoder-based language model SinBERT, by fine-tuning an encoder-based Transformer [21] with 15.7 million Sinhala sentences.
@@ III Model Selection
In the initial phase of our research, we aimed to identify the most suitable open-source multiLLM, focusing on LRLs. The first round of was carried out referring to empirical research that compared open-source LLMs. This resulted in the following models: Llama-3-8B [1], MaLA-500 [8], Aya-101 [9], PolyLM [22], XGLM, [23] Falcon [24], BLOOMZ [25], Phi3 [26], Mistral [12], OPT [27], Aya-23 [9], and BigTrans [28]. These models are listed in Table I. Based on criteria including multilingual capability, parameter size (fewer than 8 billion parameters) [^2] Model size has a direct relationship with the memory requirement in using an LLM., Sinhala language representation, and performance metrics, we selected Llama-3-8B [1] for further experimentation.
@@
Model Parameters
Official
Sinhala Support
Number of
languages
@@ IV-A Pre-training Data
Pre-training of an LLM has to be done with a substantially large dataset, in order to guarantee it learns the language-specific information. Therefore, we created a dataset by combining MADLAD-400 [33] and CulturaX [34] for the Sinhala language [^3] The dataset created by Dhananjaya et al. [7] is no longer accessible. According to Ranathunga et al. [35], this is a common problem in LRL domain.. MADLAD-400 is a manually audited, general-domain 3 Trillion token monolingual dataset based on CommonCrawl [36], spanning 419 languages. CulturaX is a multilingual dataset with 6.3 Trillion tokens in 167 languages. Although both have been cleaned, there were some entries that contained non-Sinhala words. Because of that, heuristics-based filtering [37] was applied to remove such noisy sentences. These include rules such as minimum sentence length, non-Sinhala sentence removal, exclusion of sentences with irregular punctuation, and elimination of URLs and numeric prefixes.For the pre-training, we combined both datasets and removed duplicates. Our final dataset contains 10,730,154 (10.7M) Sinhala sentences, comprising 303,958,959 (303.96M) tokens. This dataset is publicly released [^4] https://huggingface.co/datasets/polyglots/MADLAD_CulturaX_cleaned.
@@ V Continual Pretraining to Build SinLlama
We used vocabulary extension and continuous pre-training (see Section II-C) on Llama-3-8B to create SinLlama. We trained a tokenizer using the pre-training corpus using tiktoken [^5] https://github.com/openai/tiktoken, and merged this tokenizer with the Llama 3 tokenizer. We used this extended tokenizer and performed continual pretraining using the codebase from Chinese-Llama [40]. The only modification made to the original Chinese-Llama hyperparameters was the reduction of the block_size from 1024 to 512666Larger block sizes cause memory allocation issues on the GPUs we used, preventing successful model loading..
@@ VI-A Prompt Selection
For both inference and supervised fine-tuning, it was essential to select the most effective prompt templates from the available datasets. We adopted the prompt format introduced by the Alpaca Instruct Template [41] as in Fig. 1. To identify the most suitable prompts for each classification task, we referred to the work by Ahuja et al. [13], which provided well-designed prompts for several tasks.
@@ VIII Conclusion
Acknowledgment
We thank CSIR - Central Scientific Instruments Organization, India, and Emojot (Pvt) Ltd for providing the computing resources.
@@ A-A Writing-style-classification
Feature columns in the dataset are “comments”, “labels”.
@@ A-B Sentiment-tagger
The original dataset consist of XML tags. After the processing we created a dataset with the following features. Feature columns in the dataset are “context”, “text”, “sentiment”.
@@ A-C Sinhala-News-Category-classification
Feature columns in the dataset are “comments”, “labels”.